Ticket Number: HEA-NOC/20070827-7 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: ESBT
Ticket Opened: 20070827 17:26 UTC+1 Problem Start: 20070916 06:00 UTC+1
Ticket Closed: 20070917 17:08 UTC+1 Problem End: 20070916 08:00 UTC+1
Site/line: GTI, LIT, Bluenet CW - GAL, Bluenet CW - LIM, Bluenet
ESBT require a 2 hour outage on the NTFON Fibre cable between Galway
(Ballybane) and Drumline. This will also affect unprotected DWDM
wavelengths from Galway to Killonan (Limerick). For Health and Safety
reasons this work must be carried out in daylight hours.
Loss of NBE resilience:
Galway <-> Limerick
Galway <-> Citywest
Limerick <-> Citywest
20070827 15:20 CS Opening ticket.
20070913 09:31 BB BB Simultaneous work being carried out on DKIT's power
systems (detailed in 20070827-7) will result in likely loss of service for
GTI and loss of resilience for LIT during this two hour window.
20070917 17:08 GAR RFO received from ESBT.
Reason for Outage Report
Issued by ESB Telecoms Operations Centre on 11th September 2007
OP INC 392573
Service affected: ESB Telecoms managed DWDM and SDH services on the NTFON
Finglas to Mullaharlin, Dundalk.
There was a failure of all some cores of the NTFON fibre cable between
Finglas and Mullaharlin on Thur 6th September at 20:24. Investigation
revealed that cores 33-48 had been damaged by shot gun pellets. Engineers
were dispatched to move affected services to working cores. Services were
restored at 23:26
1.Confirm solved status of Problem.
The fault incident was recorded as solved at 09:15 on 8th Sept 2007.
Outage duration 18 hours.
A full repair and restoration was of services was carried out.
2.Establish and review the Sequence of Events
Thur 06/09/07 20:24 ESB Telecoms On Call Engineer (OCE) received alarms
Loss of signal from MUX kit at Finglas and Mullaharlin.
20:40 ESBT On Call support initial investigation suspects a fibre break
and calls the fibre support team.
21:15 Fibre support engineers dispatched.
Damage to fibre confirmed on tube 4.
23:26 Services moved to undamaged cores.
3.Review customer communication.
Customers where updated at regular relevant intervals by email and phone.
4.Identify issues to be addressed.
Full repair will be carried out mid October. Outage notice from ESB TOC to
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20070827-7
HEAnet Limited. Registered in Ireland, No. 275301.